OpenStudy (cjn1130):

The measures of two complementary angles are in the ratio 2:3. What is the measure of the smaller angle? I don't know how to start this problem.

OpenStudy (dumbcow):

complementary angles must add to 90 degrees 2 angles are: x , 90-x ratio means fraction \[\rightarrow \frac{x}{90-x} = \frac{2}{3}\] solve for x by cross-multiplying ---> 3x = 2(90-x) ...

OpenStudy (cjn1130):

okay thanks so much! 3x=180-2x 5x=180 x=36 so the answer is 36 right?